# Magnetocalorics of singlet ground state induced moment magnets

### v1

@@ VII Discussion of numerical results
The basis for understanding the physical effects in the two singlet model is the determination of the induced oder parameters given by the solutions of the selfconsistent solutions of Eq. (10). Two examples $(\xi_{s}=1.2)$ for FM (temperature dependence) and AFM (field dependence) are shown in Fig. 1. For the former the zero-field FM transition immediately turns into a crossover behavior at finite field since the appearance of a magnetization $\langle S_{x}\rangle$ is no longer a symmetry breaking in external fields. The same is true for the renormalized singlet splitting $\hat{\Delta}_{T}$. The AF case shown in (b) demonstrates the two sublattice order parameters $\langle S^{A,B}_{x}\rangle$ as well as associated staggered and homogeneous magnetization as function of field up to the critical one $h_{c}(T)$. The corresponding field dependence of renormalized splittings $\Delta^{\lambda}_{T}(h)$ for A,B sublattices is non-monotonic as function of field, in particular for B sublattice which is originally polarized opposite to the external field. This is quite distinct from the FM case and has consequences for the magnetocaloric effects as discussed below.
@@
The homogeneous susceptibility anomaly is a primary method to characterize H-T phase diagrams by tracing their anomalies. In Fig. 3 we present the results for both cases. Since only $J_{z}$ operator has nonzero matrix elements between the singlet only a longitudinal (zz) susceptibility has to be considered. For FM the singularity at $T_{m}$ that signifies the order is rapidly suppressed in an external field. As shown below the suppression occurs in parallel with that of the specific heat anomaly at the transition. In the AFM the q= 0 susceptibility stays finite at the transition with a cusp anomaly appearing that shifts with $T_{m}(h)$ to lower values (c.f. Fig. 2(a)) until it vanishes at the critical field $h_{c}$. For higher field a plateau due to the pure vanVleck term is present which is shifted downward because of the increasing singlet splitting. An important point is the depression $\delta\chi_{AFM}(h)$ (Eq. (28)) from peak height at $T_{m}(h)$ to zero temperature. In degenerate ground state moments the longitudinal susceptibilty drops to zero for the fully polarized AFM at T=0. This is not the case in an induced moment magnet because the saturation moment (in units of $g_{J}\mu_{B}$) is $\frac{m_{s}}{2\xi_{s}}(\xi_{s}^{2}-1)^{\frac{1}{2}}$ depends on $\xi_{s}$ and is smaller than the maximum value $\frac{1}{2}m_{s}$, reaching zero at the QCP $\xi_{s}$=1. Therefore it still can be polarized by a probe field and hence the longitudinal susceptibility is nonvanishing at $T=0$. At zero field the depression $\delta\chi_{AFM}(0)$ is given by Eq. (28) and only on approaching the degenerate case for $\xi_{s}\gg 1$ it reaches maximum value one, i.e. $\chi(0)$ vanishes. For fixed $\xi_{s}$ and increasing field the opposite is observed: The depression becomes less and vanishes at the critical field as summarized in the inset of Fig. 3. This variable depression $\delta\chi_{AFM}(h;\xi_{s})$ of the longitudinal susceptibility, depending on field as well as control parameter is characteristic for the induced moment magnets and might be used to quantify $\xi_{s}$ from experiment in the AFM case.
@@
The specific heat evolution of the singlet-singlet magnet as function of field and temperature is shown in Figs. 4 and 5 for AFM and FM cases, respectively. In both figures the dashed black curve corresponds to the background Schottky anomaly $C_{V}^{0}(T)$ resulting from the noninteracting split singlet-singlet CEF states and the full black curve corresponds to the interacting zero field case with $\xi_{s}=1.2$. It is valid for both FM and AFM transition at zero field. For this control parameter the specific heat jump of the zero field transition $(T_{m})$ sits right on top of the background Schottky anomaly peak $(T_{max})$. Then for AFM case in Fig. 4 below $T_{m}$ the entropy release is due to the ordering as well as due to the background depopulation of the excited singlet as further discussed below.
@@
For finite field the FM transition in Fig. 5 at $T_{m}$ rapidly turns into a crossover and the transition peak merges with the background. On the other hand for the AFM case the ordering at $T_{m}(h)$ prevails at finite field but shifts to lower temperature (see also Fig. 2(a)) with the specific heat jump sliding down on the left flank of the background Schottky anomaly and being continuously reduced in size. The $\delta C_{V}(T_{m}(h))$ jump normalized to the maximum of $C_{V}^{0}(T_{max})=0.44k_{B}/\text{site}$ is shown in the inset as function of field up to the critical $h_{c}$ for this control parameter where it vanishes.
As a specific example of the competing presence of Schottky anomaly and ordering specific heat jump we have discussed the case of ferromagnetic PrI3 in Sec. V as presented in Fig. 6 where we showed that from the relative position of the two peaks one may estimate the control parameter as $\xi_{s}\approx 1.08$ in this case, which is already rather close to the QCP of $\xi^{c}_{s}=1$ below which magnetic order is suppressed. We stress that we do not intend this as a quantitative comparison of temperature dependences because PrI3  has a linear low temperature itinerant specific heat contribution (subtracted in Fig. 6) and also presumably contributions from higher CEF states at elevated temperatures (the entropy exceeds $ln2$ there), both effects are not described by our two-singlet model. But the comparison in Fig. 6 demonstrates a similarity of shapes for $C_{V}/T$: A jump due to FM order sitting right on top of the Schottky peak. The essential point we want to make is that the relative position $T_{max}$ of low temperature Schottky background maximum and ordering temperature Tm gives information on the control parameter and the associated distance to the FM QCP of the induced moment magnetism.
@@
The magnetocaloric properties of the singlet-singlet model are also presented by the contour plots of $C_{V}$ in Fig. 7 in the field-temperature plane. The comparison of FM and AFM cases shows their distinct appearance since the magnetic phase transition survives in the latter up to the critical field while in the former it is rapidly suppressed and turned into a crossover. The latter is seen in Fig. 7(a),(c) for two different control parameters $\xi_{s}$. For finite field the polarization $\langle S_{x}\rangle$ and hence the singlet effective splitting $\hat{\Delta}_{T}(h)$ increase rapidly with field shifting the Schottky anomaly to higher temperature. The AFM case shows more intricate behavior in Fig. 7(b),(d): The contour of the phase boundary $T_{m}(h)$ is clearly visible but becomes blurred at lower $T_{m}$ on approaching $h_{c}$ because the specific heat jump is strongly suppressed in this region (see inset of Fig. 4). For $\xi=1.1$ the phase boundary $T_{m}(h)$ and broad Schottky maximum at $T_{max}(h)$ are separated while for $\xi_{s}=1.2$ they coincide. The latter shows now much less field dependence because there are two singlet excitation energies $\hat{\Delta}^{\lambda}_{T}(h)$ at the two sublattices $\lambda=A,B$. One of then (A) increases and the other (B) decreases (because the polarization is opposite to the external field, see Fig.1(b)). Then the thermal averaging leads to a weak field dependence of the background Schottky anomaly in this case for fields up to the critical field $h_{c}$.
@@
We emphasize that the two coexisting channels of entropy release, depopulation and ordering are characteristic for the induced moment magnetism. In a degenerate ground state magnet only the latter is active. We show this explicitly in Fig. 8 which presents the comparison of total entropy $S(T)$ and background single-site entropy $S_{0}(T)$ in the upper panel. The right panel presents the relative entropy from ordering compared to the one from depopulation, ie., the ratio $(S_{0}(T)-S(T))/(\ln 2-S_{0}(T))$ where $\ln 2$ is the high temperature limit of S(T). It demonstrates that it increases strongly with $\xi_{s}$ when the release from ordering becomes dominant and finally exclusive in the limit $\xi_{s}\gg 1$ where the level splitting will be irrelevant.
@@
The field dependence of specific heat also provides the basis for the adiabatic magnetocaloric cooling rate $\Gamma_{mc}(T,h)$ as given by Eq. (40) and shown for the FM and AFM case in Fig. 9. In the former when there is no phase transition at finite field it behaves smoothly, increasing a lower field and decreasing at lower temperature mostly due to the behavior of the temperature derivative of the magnetization in Fig. 9. For $h\rightarrow 0$ it diverges when $T\rightarrow T_{m}^{-}$ on approaching the FM transition. This is because $M_{x}^{0}$ becomes the FM order parameter in that limit with a diverging temperature derivative. On the other hand for $T>T_{m}$ when $M_{x}^{0}$ is no longer singular for small fields $\Gamma_{mc}$ will drop to zero. For larger fields it will fall off $\sim 1/h$ in agreement with the noninteracting case (Appendix A).
@@
Instead of sweeping the field one may apply rapid adiabatic pressure change leading to the analogous barocaloric cooling rate $\Gamma_{bc}(T,p)$ at a constant applied field (Eq. (41)). It is shown in Fig. 10 as function of pressure and for different fields. For zero field the FM and AFM pressure dependence of $\Gamma_{bc}$ is identical. The exhibit a sign change and a jump at the critical pressure above which magnetic order vanishes (due to the decrease of $\xi_{s}(p)$). For finite fields they are largely different. While for the FM the barocaloric coefficient develops a smooth pressure dependence in the AFM ordered regime the step prevails and moves to lower critical pressures as field values increase up to the critical one where it vanishes. In both cases the step at the critical pressure moves to lower or higher values when temperature increases or decreases, respectively.
[…]
The more interesting AFM case is presented in the lower panels. Fig. 11(c) shows the magnetic exciton dispersion (Eq. (VI)) of the AFM right at the transition temperature where the lower mode $\omega_{0}({\bf Q})$ becomes soft at the AFM ordering vector (R-point). For larger field but inside the AFM phase two modes appear due to the downfolding to AFBZ and gapping and hybridization at the new AFBZ zone boundaries which is due to the fact that $\hat{\Delta}_{T}^{A}\neq\hat{\Delta}_{T}^{B}$ for the two sublattices. The temperature dependence of the lower mode for fields increasing from zero to $h_{c}$ is shown in Fig. 12. The softening in the para phase above $T_{m}(h)$ is clearly seen but the soft mode appears only at $h=0,h_{c}$. For fields inside the interval $(0,h_{c})$ the soft mode is arrested due to the subtle effect of the pseudo-Curie terms on the the transition temperature as explained before in Sec. VI. The resulting arc-like field dependence of $\omega_{-}({\bf Q},T_{m}(h),h)$ visible in Fig. 12 is even more clearly seen in Fig. 2(a) based on the analytical formula in Eq. (VI).
VIII Summary and Conclusion
In this work we have studied primarily the magnetocalorics and field-dependence of magnetic exciton modes of a simple singlet-singlet model. It describes schematically the magnetism in numerous 4f and 5f compounds with integer occupied f shells (non-Kramers ions) in a radically simplified manner: All excited CEF states are assumed to be at sufficiently high energies as compared to the transition temperature and the lowest CEF singlet-singlet splitting. Since both singlet states are nonmagnetic the magnetism in this model can only occur by a spontaneuos admixture of ground- and excited states generating an induced moment.
The specific heat appearance is connected with the possible low temperature entropy release shows two contributions: Firstly from the upper level depopulation leading to the Schottky-type background anomaly with a maximum at $T_{max}$ and secondly a specific heat jump at the critical temperature $T_{m}$ for induced order. Their relative position is determined by the control parameter. For $\xi_{s}\approx 1.2$ the two anomalies are on top of each other and both give a significant contribution to the entropy release. The recently found induced moment singlet-singlet FM compound PrI3 is rather close to that situation.
In external field only the AFM transition can survive whereas the FM changes to crossover behavior. In the former case the specific heat jump at $T_{m}(h)$ is shifted to lower temperature sliding on the lower flank of the almost field independent Schottky anomaly and reducing is size until it vanishes at the critical field. This means that for a control parameter where the zero field $T_{m}$ and $T_{max}$ are close they will rapidly be suppressed in an applied field. This should be a distinct characteristic of induced moment AFM. On the other hand for the FM the specific heat jump at zero field is quickly absorbed in the Schottky anomaly that shifts to higher temperatures with increasing field.
The rather different magnetocaloric characteristics of induced singlet FM and AFM are also reflected in the adiabatic cooling rate. For FM it is continuously increasing with lowering field or increasing temperature. On the other hand for the AFM it changes sign at $T_{m}(h)$ and is strongly reduced below reaching zero at low fields. And a similar distinction is observed in the barocaloric coefficient.
@@ VIII Summary and Conclusion
The critical properties of the induced moment magnets are also reflected in the excitation spectrum, i.e. the magnetic exciton dispersions and their field and temperature dependence within and outside the ordered region. In FM and AFM zero field case a soft mode appears at the respective ordering wave vectors ${\bf q}=0,{\bf Q}$ at $T_{m}$ and a re-hardening of the mode occurs below the transition. In the AFM case a surprising field dependence is observed: When moving along the phase boundary $T_{m}(h)$ a true soft mode appears only at the endpoints $h=0,h_{c}$, in between an arrested soft mode with finite frequency is observed. This was shown to be due to induced pseudo-Curie term in the static susceptibility that lead to a $T_{m}(h)$ larger than the temperature where complete softening could be achieved.
@@
The numerous specific magnetocaloric and field- dependent magnetic exciton dispersive characteristics for the singlet induced moment magnets that have been derived in this work should be helpful to characterize localized f-electron magnets that may be approximately correspond to this class, depending on the concrete CEF level scheme of the compound.

### v2

@@ VII Discussion of numerical results
The basis for understanding the physical effects in the two singlet model is the determination of the induced oder parameters given by the solutions of the selfconsistent solutions of Eq. (10). Two examples $(\xi_{s}=1.2)$ for FM (temperature dependence) and AFM (field dependence) are shown in Fig. 1. For the former the zero-field FM transition immediately turns into a crossover behaviour at finite field since the appearance of a magnetisation $\langle S_{x}\rangle$ is no longer a symmetry breaking in external fields. The same is true for the renormalised singlet splitting $\hat{\Delta}_{T}$. The AF case shown in (b) demonstrates the two sublattice order parameters $\langle S_{x}\rangle_{A,B}$ as well as associated staggered and homogeneous magnetisation as function of field up to the critical one $h_{c}(T)$. The corresponding field dependence of renormalised splittings $\Delta^{\lambda}_{T}(h)$ for A,B sublattices is non-monotonic as function of field, in particular for B sublattice which is originally polarised opposite to the external field. This is quite distinct from the FM case and has consequences for the magnetocaloric effects as discussed below.
@@
The homogeneous susceptibility anomaly is a primary method to characterise H-T phase diagrams by tracing their anomalies. In Fig. 3 we present the results for both cases. Since only $J_{z}$ operator has nonzero matrix elements between the singlet only a longitudinal (zz) susceptibility has to be considered. For FM the singularity at $T_{m}$ that signifies the order is rapidly suppressed in an external field. As shown below the suppression occurs in parallel with that of the specific heat anomaly at the transition. In the AFM the q= 0 susceptibility stays finite at the transition with a cusp anomaly appearing that shifts with $T_{m}(h)$ to lower values (c.f. Fig. 2(a)) until it vanishes at the critical field $h_{c}$. For higher field a plateau due to the pure vanVleck term is present which is shifted downward because of the increasing singlet splitting. An important point is the depression $\delta\chi_{AFM}(h)$ (Eq. (28)) from peak height at $T_{m}(h)$ to zero temperature. In degenerate ground state moments the longitudinal susceptibility drops to zero for the fully polarised AFM at T=0. This is not the case in an induced moment magnet because the saturation moment (in units of $g_{J}\mu_{B}$) is $\frac{m_{s}}{2\xi_{s}}(\xi_{s}^{2}-1)^{\frac{1}{2}}$ which depends on $\xi_{s}$ and is smaller than the maximum value $\frac{1}{2}m_{s}$, reaching zero at the QCP $\xi_{s}$=1. Therefore it still can be polarised by a probe field and hence the longitudinal susceptibility is nonvanishing at $T=0$. At zero field the depression $\delta\chi_{AFM}(0)$ is given by Eq. (28) and only on approaching the degenerate case for $\xi_{s}\gg 1$ it reaches maximum value one, i.e. $\chi(0)$ vanishes. For fixed $\xi_{s}$ and increasing field the opposite is observed: The depression becomes less and vanishes at the critical field as summarised in the inset of Fig. 3. This variable depression $\delta\chi_{AFM}(h;\xi_{s})$ of the longitudinal susceptibility which depends on field as well as control parameter is characteristic for the induced moment magnets and might be used to quantify $\xi_{s}$ from experiment in the AFM case.
Before discussing the specific heat it is useful to consider an important symmetry aspect of the induced magnetic order. For FM only time reversal symmetry is broken while for AFM both time reversal and translational symmetry are broken (although the product is still conserved). Application of an external field removes time reversal symmetry and then in the FM there is no more symmetry breaking because the homogeneous magnetisation $M_{x}^{0}$ (Eq.(11)), which is the order parameter at zero field is now already present above $T_{m}(h=0)$. Therefore at finite field no phase transition but only a gradual crossover to a more polarised state will appear around and below this temperature. This leads to the rapid suppression of the zero-field specific heat jump in finite field. On the other hand for AFM the translational symmetry breaking due to the two sublattices is still present in finite field, characterised by the staggered order parameter $M_{x}^{s}$ (Eq.(11)). Consequently the induced AFM phase transition survives up to the critical field as long as $M_{x}^{s}$ is finite (Fig. 1(b)). These different symmetry properties of the two types of order will be reflected in their strikingly distinct field dependence of the specific heat.
@@
The specific heat evolution of the singlet-singlet magnet as function of field and temperature is shown in Figs. 4(a,b) for AFM and FM, respectively. In both figures the dashed black curve corresponds to the background Schottky anomaly $C_{V}^{0}(T)$ resulting from the noninteracting split singlet-singlet CEF states and the full black curve corresponds to the interacting zero field case with $\xi_{s}=1.2$. It is valid for both FM and AFM transition at zero field. For this control parameter the specific heat jump of the zero field transition $(T_{m})$ sits right on top of the background Schottky anomaly peak $(T_{max})$. Then for AFM case in Fig. 4(a) below $T_{m}$ the entropy release is due to the ordering as well as due to the background depopulation of the excited singlet as further discussed below.
@@
For finite field the FM transition in Fig. 4(b) at $T_{m}$ rapidly turns into a crossover and the transition peak merges with the background. On the other hand for the AFM case the ordering at $T_{m}(h)$ prevails at finite field but shifts to lower temperature (see also Fig. 2(a)) with the specific heat jump sliding down on the left flank of the background Schottky anomaly and being continuously reduced in size. The $\delta C_{V}(T_{m}(h))$ jump normalised to the maximum of $C_{V}^{0}(T_{max})=0.44k_{B}/\text{site}$ is shown in the inset as function of field up to the critical $h_{c}$ for this control parameter where it vanishes.
As a specific example of the competing presence of Schottky anomaly and ordering specific heat jump we have discussed the case of ferromagnetic PrIr3 in Sec. V.1 as presented in Fig. 5 where we showed that from the relative position of the two peaks one may estimate the control parameter as $\xi_{s}\approx 1.08$ in this case, which is already rather close to the QCP of $\xi^{c}_{s}=1$ below which magnetic order is suppressed. The field-dependence of the magnetic specific heat of PrIr3 which should look qualitatively similar to Fig. 4(b) has not yet been determined experimentally.
@@
The magnetocaloric properties of the singlet-singlet model are also presented by the contour plots of $C_{V}$ in Fig. 6 in the field-temperature plane. The comparison of FM and AFM cases shows their distinct appearance since the magnetic phase transition survives in the latter up to the critical field while in the former it is rapidly suppressed and turned into a crossover. The latter is seen in Fig. 6(a),(c) for two different control parameters $\xi_{s}$. For finite field the polarisation $\langle S_{x}\rangle$ and hence the singlet effective splitting $\hat{\Delta}_{T}(h)$ increase rapidly with field shifting the Schottky anomaly to higher temperature. The AFM case shows more intricate behaviour in Fig. 6(b),(d): The contour of the phase boundary $T_{m}(h)$ is clearly visible but becomes blurred at lower $T_{m}$ on approaching $h_{c}$ because the specific heat jump is strongly suppressed in this region (see inset of Fig. 4(a)). For $\xi=1.1$ the phase boundary $T_{m}(h)$ and broad Schottky maximum at $T_{max}(h)$ are separated while for $\xi_{s}=1.2$ they coincide. The latter shows now much less field dependence because there are two singlet excitation energies $\hat{\Delta}^{\lambda}_{T}(h)$ at the two sublattices $\lambda=A,B$. One of then (A) increases and the other (B) decreases (because the polarisation is opposite to the external field, see Fig.1(b)). Then the thermal averaging leads to a weak field dependence of the background Schottky anomaly in this case for fields up to the critical field $h_{c}$.
@@
We emphasise that the two coexisting channels of entropy release, depopulation and ordering are characteristic for the induced moment magnetism. In a degenerate ground state magnet only the latter is active. We show this explicitly in Fig. 7 which presents the comparison of total entropy $S(T)$ and background single-site entropy $S_{0}(T)$ in the upper panel. The right panel presents the relative entropy from ordering compared to the one from depopulation, ie., the ratio $(S_{0}(T)-S(T))/(\ln 2-S_{0}(T))$ where $\ln 2$ is the high temperature limit of S(T). It demonstrates that it increases strongly with $\xi_{s}$ when the release from ordering becomes dominant and finally exclusive in the limit $\xi_{s}\gg 1$ where the level splitting will be irrelevant.
@@
The field dependence of specific heat also provides the basis for the adiabatic magnetocaloric cooling rate $\Gamma_{mc}(T,h)$ as given by Eq. (40) and shown for the FM and AFM case in Fig. 8. In the former, when there is no phase transition at finite field it behaves smoothly, increasing at lower field and decreasing at lower temperature mostly due to the behaviour of the temperature derivative of the magnetisation in Fig. 8. For $h\rightarrow 0$ it diverges when $T\rightarrow T_{m}^{-}$ on approaching the FM transition. This is because $M_{x}^{0}$ becomes the FM order parameter in that limit with a diverging temperature derivative. On the other hand for $T>T_{m}$ when $M_{x}^{0}$ is no longer singular for small fields $\Gamma_{mc}$ will drop to zero. For larger fields it will fall off $\sim 1/h$ in agreement with the noninteracting case (Appendix A).
For the AFM case pronounced anomalies and a sign change of $\Gamma_{mc}$ are seen at the critical field $h_{c}$. In the ordered phase as well in the para phase ($T>T_{m}$) it approaches zero at small fields since the temperature derivative of $M_{x}^{0}$ always tends to zero in that limit for the AFM. For fields $h\gg h_{c}$ and much larger than in the FM case $\Gamma_{mc}$ will achieve a maximum and then fall off like $1/h$ as in the noninteracting limit. The maximum field in Fig. 8(b) is still below that range. The integrated adiabatic temperature change when decreasing the field is negative at all fields for FM while it is negative above $h_{c}$ and slightly positive below for the AF case.
Instead of sweeping the field one may apply rapid adiabatic pressure change leading to the analogous barocaloric cooling rate $\Gamma_{bc}(p)$ at a constant applied field (Eq. (41)). It is shown in Fig. 9 as function of pressure for different fields and fixed temperature. For zero field the FM and AFM pressure dependence of $\Gamma_{bc}$ is identical. The exhibit a sign change and a jump at the critical pressure above which magnetic order vanishes (due to the decrease of $\xi_{s}(p)$). For finite fields they are largely different. While for the FM the barocaloric coefficient develops a smooth pressure dependence in the AFM ordered regime the step prevails and moves to lower critical pressures as field values increase up to the critical one where it vanishes. If the field is kept constant in both FM and AFM cases but temperature is increased, then in each of the curves in Fig. 9 the step anomaly (broadened for FM) is shifted to lower pressure, but the shape of the $\Gamma_{bc}(p)$ curve is qualitatively preserved.
Finally we discuss the field and temperature dependence of the singlet-singlet magnetic exciton spectrum as presented in the panels in Fig. 10. For FM as well as AFM mode presentation we use the paramagnetic simple cubic BZ. It is an extended zone scheme for the AFM. In this case the $\Gamma$- and R-points are equivalent points in the AFM phase connected by ${\bf Q}=(111)$ ordering vector and correspond both to centers of the fcc AFBZ. Their midpoint L($\frac{1}{2},\frac{1}{2},\frac{1}{2})$) is the AFBZ zone boundary point.
The more interesting AFM case is presented in the lower panels. Fig. 10(c) shows the magnetic exciton dispersion (Eq. (VI)) of the AFM right at the transition temperature where the lower mode $\omega_{0}({\bf Q})$ becomes soft at the AFM ordering vector (R-point). For larger field but inside the AFM phase two modes appear due to the downfolding to AFBZ accompanied by gapping and hybridization at the new AFBZ zone boundaries. This is due to the fact that $\hat{\Delta}_{T}^{A}\neq\hat{\Delta}_{T}^{B}$ for the two sublattices. The temperature dependence of the lower mode for fields increasing from zero to $h_{c}$ is shown in Fig. 11. The softening in the para phase above $T_{m}(h)$ is clearly seen but the soft mode appears only at $h=0,h_{c}$. For fields inside the interval $(0,h_{c})$ the soft mode is arrested due to the subtle effect of the pseudo-Curie terms on the the transition temperature as explained before in Sec. VI. The resulting arc-like field dependence of $\omega_{-}({\bf Q},T_{m}(h),h)$ visible in Fig. 11 is even more clearly seen in Fig. 2(a) based on the analytical formula in Eq. (VI).
VIII Summary and Conclusion
In this work we have studied primarily the magnetocalorics and field-dependence of magnetic exciton modes of a simple singlet-singlet model. It describes schematically the magnetism in numerous 4f and 5f compounds with integer occupied f shells (non-Kramers ions) in a radically simplified manner: All excited CEF states are assumed to be at sufficiently high energies as compared to the transition temperature and the lowest CEF singlet-singlet splitting. Since both singlet states are nonmagnetic the magnetism in this model can only occur by a spontaneous admixture of ground- and excited states generating an induced moment.
The specific heat appearance is connected with the possible low temperature entropy release shows two contributions: Firstly from the upper level depopulation leading to the Schottky-type background anomaly with a maximum at $T_{max}$ and secondly a specific heat jump at the critical temperature $T_{m}$ for induced order. Their relative position is determined by the control parameter. For $\xi_{s}\approx 1.2$ the two anomalies are on top of each other and both give a significant contribution to the entropy release. The recently found induced moment singlet-singlet FM compound PrIr3 is rather close to that situation.
In external field only the AFM transition can survive whereas the FM changes to crossover behaviour. In the former case the specific heat jump at $T_{m}(h)$ is shifted to lower temperature sliding on the lower flank of the almost field independent Schottky anomaly and reducing is size until it vanishes at the critical field. This means that for a control parameter where the zero field $T_{m}$ and $T_{max}$ are close they will rapidly be suppressed in an applied field. This should be a distinct characteristic of induced moment AFM. On the other hand for the FM the specific heat jump at zero field is quickly absorbed in the Schottky anomaly that shifts to higher temperatures with increasing field.
@@ VIII Summary and Conclusion
The critical properties of the induced moment magnets are also reflected in the excitation spectrum, i.e., the magnetic exciton dispersions and their field and temperature dependence within and outside the ordered region. In FM and AFM zero field case a soft mode appears at the respective ordering wave vectors ${\bf q}=0,{\bf Q}$ at $T_{m}$ and a re-hardening of the mode occurs below the transition. In the AFM case a surprising field dependence is observed: When moving along the phase boundary $T_{m}(h)$, a true soft mode is found only at the endpoints $h=0,h_{c}$, in between an arrested soft mode with finite frequency is observed. This was shown to be due to induced pseudo-Curie term in the static susceptibility that lead to a $T_{m}(h)$ larger than the temperature where complete softening could be achieved.
@@
The numerous specific magnetocaloric and field- dependent magnetic exciton dispersive characteristics for the singlet induced moment magnets that have been derived in this work should be helpful to characterise localised f-electron magnets that may be approximately correspond to this class, depending on the concrete CEF level scheme of the compound.